# Theory on Mixture-of-Experts in Continual Learning

### v2

@@ 1 Introduction
Continual Learning (CL) has emerged as an important paradigm in machine learning [1, 2], in which an expert aims to learn a sequence of tasks one by one over time. The expert is anticipated to leverage the knowledge gained from old tasks to facilitate learning new tasks, while simultaneously enhancing the performance of old tasks via the knowledge obtained from new ones. Given the dynamic nature of CL, one major challenge herein is known as catastrophic forgetting [3, 4], where the expert can perform poorly on (i.e., easily forget) the previous tasks when learning new tasks if data distributions change largely across tasks. This becomes a more serious issue if a single expert continues to serve an increasing number of tasks.
Recently, the sparsely-gated Mixture-of-Experts (MoE) model has achieved astonishing successes in deep learning, especially in the development of large language models (LLMs) (e.g., [5, 6, 7, 8]). By adaptively routing different input data to one of the multiple experts through a gating network [9, 10], different experts in the MoE model will be specialized to grasp different knowledge in the data. Thus inspired, there have emerged several attempts to leverage MoE in mitigating the forgetting issue in CL (e.g., [11, 12, 13, 14, 15]) by training each expert to handle a particular set of tasks. However, these studies have primarily focused on experimental investigations, whereas the theoretical understanding of MoE and its impact on the learning performance in CL is still lacking. In this paper, we aim to fill this gap by providing the first explicit theoretical results to comprehensively understand MoE in CL.
To this end, we study the sparsely-gated MoE model with $M$ experts in CL through the lens of overparameterized linear regression tasks [16, 17, 18]. In our setting of CL, one learning task arrives in each round and its dataset is generated with the ground truth randomly drawn from a shared pool encompassing $N$ unknown linear models. Subsequently, the data is fed into a parameterized gating network, guided by which a softmax-based router will route the task to one of the $M$ experts for model training. The trained MoE model then further updates the gating network by gradient descent (GD). After that, a new task arrives and the above process repeats until the end of CL. It is worth noting that analyzing linear models is an important first step towards understanding the performance of deep neural networks (DNNs), as shown in many recent studies (e.g., [19, 20, 21]).
@@
We provide explicit expressions of the expected forgetting and generalization error to characterize the benefit of MoE on the performance of CL. 1) Compared to the single expert case ($M=1$), where tasks are learned by a single diverged model, the MoE model with diversified experts significantly enhances the learning performance, especially with large changes in data distributions across tasks. 2) Regardless of whether there are more experts ($M>N$) or fewer experts ($M<N$), both forgetting and generalization error converge to a small constant. This occurs because the router consistently selects the right expert for each task after the MoE model converges, efficiently minimizing model errors in the long term. 3) In MoE, adding more experts requires additional rounds for exploration before convergence, which does not necessarily enhance the learning performance.
@@ 2 Related work
Continual learning. In the past decade, various empirical approaches have been proposed to tackle catastrophic forgetting in CL, which generally fall into three categories: 1) Regularization-based approaches (e.g., [4, 23, 24, 25]), which introduce explicit regularization terms on key model parameters trained by previous tasks to balance old and new tasks. 2) Parameter-isolation-based approaches (e.g., [26, 27, 28, 29]), which isolate parameters associated with different tasks to prevent interference between parameters. 3) Memory-based approaches (e.g., [30, 31, 32, 33, 34, 35]), which store data or gradient information from old tasks and replay them during training of new tasks.
@@
On the other hand, theoretical studies on CL are very limited. Among them, [36] and [37] introduce NTK overlap matrix to measure the task similarity and propose variants of the orthogonal gradient descent approach to address catastrophic forgetting. [38] considers a teacher-student framework to examine the impact of task similarity on learning performance. [39] proposes an ideal CL framework that can achieve no forgetting by assuming i.i.d. data distributions for all tasks. [19] provides forgetting bounds in overparameterized linear models on different task orders. Further, [20] provides explicit forms of forgetting and overall generalization error based on the testing error. These works collectively suggest that learning performance with a single expert tends to deteriorate when subsequent tasks exhibit significant diversification. In contrast, our work is the first to conduct theoretical analysis to understand the benefit of multiple experts on CL.
@@
Mixture-of-Experts model. The MoE model has been extensively studied over the years for enhancing model capacity in deep learning (e.g., [9, 40, 41, 42, 43]). Recently, it has found widespread applications in emerging fields such as LLMs (e.g., [5, 6, 7, 8]). To improve training stability and simplify the MoE structure, [10] proposes to sparsify the output of the gating network. Subsequently, [22] suggests routing each data sample to a single expert instead of multiple experts, which reduces computational costs while maintaining model quality. [21] theoretically analyzes the mechanism of MoE in deep learning under the setup of a mixture of classification problem. Note that this study focuses on a single-task setting, and hence does not analyze the dynamics involved in training a sequence of tasks in CL. In contrast, our work delves into MoE in CL, resulting in distinct training phases of the gating network compared to [21]. Furthermore, we provide explicit expressions for both forgetting and overall generalization error.
@@
MoE in CL. Recently, the MoE model has been applied to reducing catastrophic forgetting in CL ([11, 12, 13, 14, 15]). For example, [12] integrates the cooperation of multiple experts to make predictions based on a weighted sum of feature representations. [14] proposes to diverse experts by routing data with minimal distribution overlap to each expert and then combine experts’ knowledge during task predictions to enhance learning stability. Additionally, [15] applies MoE to expand the capacity of vision-language models, alleviating long-term forgetting in CL. However, these works solely focus on empirical methods, lacking theoretical analysis of how the MoE performs in CL.
@@ 3.1 CL in linear models
General setting. We consider the CL setting with $T$ training rounds. In each round $t\in[T]$, one out of $N$ tasks randomly arrives to be learned by the MoE model with $M$ experts. For each task, we consider fitting a linear model $f(\mathbf{X})=\mathbf{X}^{\top}\bm{w}$ [19, 20] with ground truth $\bm{w}\in\mathbb{R}^{d}$. Then, for the task arrival in the $t$-th training round, it corresponds to a linear regression problem, where the training dataset is denoted by $\mathcal{D}_{t}=(\mathbf{X}_{t},\mathbf{y}_{t})$. Here $\mathbf{X}_{t}\in\mathbb{R}^{d\times s_{t}}$ is the feature matrix with $s_{t}$ samples of $d$-dimensional vectors, and $\mathbf{y}_{t}\in\mathbb{R}^{s_{t}}$ is the output vector. In this study, we focus on the overparameterized regime, where $s_{t}<d$. Consequently, there exist numerous linear models that can perfectly fit the data.
Ground truth and dataset. Let $\mathcal{W}=\{\bm{w}_{1},\cdots,\bm{w}_{N}\}$ represent the collection of ground truth vectors of all $N$ tasks. For any two tasks $n,n^{\prime}\in[N]$, we assume $\|\bm{w}_{n}-\bm{w}_{n^{\prime}}\|_{\infty}=\mathcal{O}(\sigma_{0})$, where $\sigma_{0}>0$ denotes the variance. Moreover, we assume that task $n$ possesses a unique feature signal $\bm{v}_{n}\in\mathbb{R}^{d}$ with $\|\bm{v}_{n}\|_{\infty}=\mathcal{O}(1)$ (cf. [21]).
@@
In any training round $t\in[T]$, the actual ground truth $\bm{w}_{n_{t}}$ of task arrival $n_{t}$ is unknown. However, according to Definition 1, task $n_{t}$ can be classified into one of $N$ clusters based on its feature signal $\bm{v}_{n_{t}}$. Although the position of $\bm{v}_{n_{t}}$ in feature matrix $\mathbf{X}_{t}$ is not specified for each task $n_{t}$, we can address this binary classification sub-problem over $\mathbf{X}_{t}$ using a single gating network in MoE [10, 22, 21]. In this context, we aim to investigate whether the MoE model can enhance the learning performance in CL. For ease of exposition, we assume $s_{t}=s$ for all $t\in[T]$ in this paper. Then we will introduce the MoE model in the following subsections.
@@ 3.2 Structure of the MoE model
In each round $t$, as depicted in Figure 1, for task $n_{t}$, the router selects the expert with the maximum gate output $h_{m}(\mathbf{X}_{t},\bm{\theta}_{t}^{(m)})$, denoted as $m_{t}$, from the $M$ experts. In practice, to encourage exploration across experts and stabilize MoE training, we add perturbations to the router ([10, 22, 21]). Specifically, task $n_{t}$ will be routed to the expert that satisfies
@@
where $r_{t}^{(m)}$ for any $m\in[M]$ is drawn independently from the uniform distribution $\text{Unif}[0,\lambda]$. We analyze in Appendix B that routing strategy Eq. 1 ensures continuous and stable transitions for tasks. Additionally, the router calculates the softmaxed gate outputs, derived by
@@ 3.3 Training of the MoE model with key designs
Since we focus on the overparameterized regime, there exist infinitely many solutions that perfectly satisfy $\mathcal{L}_{t}^{tr}(\bm{w}_{t}^{(m_{t})},\mathcal{D}_{t})=0$ in Eq. 3. Among these solutions, gradient descent (GD) starting from the previous expert model $\bm{w}_{t-1}^{(m_{t})}$ at the convergent point provides a unique solution for minimizing $\mathcal{L}_{t}^{tr}(\bm{w}_{t}^{(m_{t})},\mathcal{D}_{t})$ in Eq. 3, which is determined by the following optimization problem [19, 45, 20]:
@@
Gating network parameters. After obtaining $\bm{w}_{t}^{(m_{t})}$ in Eq. 5, the MoE updates the gating network parameter from $\bm{\Theta}_{t}$ to $\bm{\Theta}_{t+1}$ using GD for the next training round. On one hand, we aim for $\bm{\theta}_{t+1}^{(m)}$ of each expert $m$ to specialize in a specific task, which helps mitigate learning loss caused by the incorrect routing of distinct tasks. On the other hand, the router needs to balance the load among all experts [22, 10, 6] to reduce the risk of model overfitting and enhance the learning performance in CL. To achieve this, we introduce the following key designs for gating network updates.
Key design I: Multi-objective training loss. First, based on the updated expert model $\bm{w}_{t}^{(m)}$ in Eq. 5, we propose the following locality loss function for updating $\mathbf{\Theta}_{t}$:
@@ 4 Theoretical results on MoE training for CL
In this section, we provide theoretical analysis for the training of expert models and the gating network in Algorithm 1, which further justifies our key designs in Section 3. Specifically, (i) we first support our key design I by proving that the expert model converges fast via updating $\bm{\Theta}_{t}$ under our designed locality loss in Eq. 6. (ii) We then show that our key design II, early termination in updating $\bm{\Theta}_{t}$ for MoE in CL, is necessary to ensure a stable convergence system state with balanced loads among all experts. For clarity, we study the case with $M>N$ in this section (where the results are labeled as $M>N$ version), and further extend the results to the case with $M<N$ in appendices.
To characterize the expert specialization, we first show that the gate output of each expert is determined by the input feature signal $\bm{v}_{n}$ of $\mathbf{X}_{t}$.
@@
The full version and the proof of Lemma 1 are given in Appendix C. According to Lemma 1, the router decides expert $m_{t}$ for task $n_{t}$ based primarily on its feature signal $\bm{v}_{n_{t}}$. Consequently, given $N$ tasks, all experts can be grouped into $N$ sets according to their specialty, i.e., their gating parameter $\bm{\theta}^{(m)}_{t}$’s to identify feature signal $\bm{v}_{n}$, where each expert set is defined as:
@@
The full version and the proof of Proposition 2 are given in Appendix F. According to Proposition 2, if the MoE keeps updating $\bm{\Theta}_{t}$ at any round $t$, the gap between gate outputs of any two experts within the same expert set converges to $\mathcal{O}(\sigma_{0}^{1.75})$ at round $t_{1}$ in Eq. 12. In contrast, for any two experts in different expert sets, their output gap is sufficiently large, i.e., $\Theta(\sigma_{0}^{0.75})$, indicating that the MoE has successfully diversified experts into different sets at round $t_{1}$ as in Eq. 11. However, unlike MoE in single-task learning that can stop training at any time after the expert models converge (e.g., [46, 6]), MoE in CL requires both the gating network and the experts to be suitably updated with the continuous arrival of new tasks. This is necessary to balance the load on each expert and maximize the system capacity utilization. However, continuing updating $\bm{\Theta}_{t}$ according to Eq. 9 will eventually reduce the output gap between any two experts to $\mathcal{O}(\sigma_{0}^{1.75})$ in Eq. 13 at training round $t_{2}$, causing the router to select wrong experts for subsequent task arrivals and incurring additional training errors.

### v3

@@ 1 Introduction
Continual Learning (CL) has emerged as an important paradigm in machine learning ([39, 50]), in which an expert aims to learn a sequence of tasks one by one over time. The expert is anticipated to leverage the knowledge gained from old tasks to facilitate learning new tasks, while simultaneously enhancing the performance of old tasks via the knowledge obtained from new ones. Given the dynamic nature of CL, one major challenge herein is known as catastrophic forgetting ([36, 23]), where the expert can perform poorly on (i.e., easily forget) the previous tasks when learning new tasks if data distributions change largely across tasks. This becomes a more serious issue if a single expert continues to serve an increasing number of tasks.
Recently, the sparsely-gated Mixture-of-Experts (MoE) model has achieved astonishing successes in deep learning, especially in the development of large language models (LLMs) (e.g., [10, 31, 32, 52]). By adaptively routing different input data to one of the multiple experts through a gating network ([11, 46]), different experts in the MoE model will be specialized to grasp different knowledge in the data. Thus inspired, there have emerged several attempts to leverage MoE in mitigating the forgetting issue in CL (e.g., [18, 49, 9, 43, 55]) by training each expert to handle a particular set of tasks. However, these studies have primarily focused on experimental investigations, whereas the theoretical understanding of MoE and its impact on the learning performance in CL is still lacking. In this paper, we aim to fill this gap by providing the first explicit theoretical results to comprehensively understand MoE in CL.
To this end, we study the sparsely-gated MoE model with $M$ experts in CL through the lens of overparameterized linear regression tasks ([48, 1, 57]). In our setting of CL, one learning task arrives in each round and its dataset is generated with the ground truth randomly drawn from a shared pool encompassing $N$ unknown linear models. Subsequently, the data is fed into a parameterized gating network, guided by which a softmax-based router will route the task to one of the $M$ experts for model training. The trained MoE model then further updates the gating network by gradient descent (GD). After that, a new task arrives and the above process repeats until the end of CL. It is worth noting that analyzing linear models is an important first step towards understanding the performance of deep neural networks (DNNs), as shown in many recent studies (e.g., [12, 34, 6]).
@@
We provide explicit expressions of the expected forgetting and generalization error to characterize the benefit of MoE on the performance of CL. 1) Compared to the single expert case ($M=1$), where tasks are learned by a single diverged model, the MoE model with diversified experts significantly enhances the learning performance, especially with large changes in data distributions across tasks. 2) Regardless of whether there are more experts ($M>N$) or fewer experts ($M<N$), both forgetting and generalization error converge to a small constant. This occurs because the router consistently selects the right expert for each task after the MoE model converges, efficiently minimizing model errors caused by switching tasks. 3) In MoE, initially adding more experts requires additional exploration rounds before convergence, which does not necessarily improve learning performance.
@@ 2 Related work
Continual learning. In the past decade, various empirical approaches have been proposed to tackle catastrophic forgetting in CL, which generally fall into three categories: 1) Regularization-based approaches (e.g., [23, 42, 16, 35]), which introduce explicit regularization terms on key model parameters trained by previous tasks to balance old and new tasks. 2) Parameter-isolation-based approaches (e.g., [5, 45, 20, 54, 24]), which isolate parameters associated with different tasks to prevent interference between parameters. 3) Memory-based approaches (e.g., [13, 21, 33, 44, 47, 15]), which store data or gradient information from old tasks and replay them during training of new tasks.
@@
On the other hand, theoretical studies on CL are very limited. Among them, [8] and [3] introduce NTK overlap matrix to measure the task similarity and propose variants of the orthogonal gradient descent approach to address catastrophic forgetting. [28] consider a teacher-student framework to examine the impact of task similarity on learning performance. [40] propose an ideal CL framework that can achieve no forgetting by assuming i.i.d. data distributions for all tasks. [12] provide forgetting bounds in overparameterized linear models on different task orders. Further, [34] provide explicit forms of forgetting and overall generalization error based on the testing error. These works collectively suggest that learning performance with a single expert tends to deteriorate when subsequent tasks exhibit significant diversification. In contrast, our work is the first to conduct theoretical analysis to understand the benefit of multiple experts on CL.
@@
Mixture-of-Experts model. The MoE model has been extensively studied over the years for enhancing model capacity in deep learning (e.g., [11, 41, 51, 58, 7, 56]). Recently, it has found widespread applications in emerging fields such as LLMs (e.g., [10, 31, 32, 52]). To improve training stability and simplify the MoE structure, [46] propose to sparsify the output of the gating network. Subsequently, [14] suggest routing each data sample to a single expert instead of multiple experts. For theoretical studies, [37] propose a maximum quasi-likelihood method for estimating MoE parameters, while [6] analyze MoE mechanisms in deep learning for single-task classification. Unlike these works, which do not address sequential task training in CL, our study focuses on MoE in CL, introducing distinct training phases for the gating network. Additionally, we derive explicit expressions for forgetting and generalization errors.
@@
MoE in CL. Recently, the MoE model has been applied to reducing catastrophic forgetting in CL ([29, 18, 49, 9, 43, 55]). For example, [29] expand the number of experts using the Bayesian nonparametric framework to address task-free CL. [43] propose to diverse experts by routing data with minimal distribution overlap to each expert and then combine experts’ knowledge during task predictions to enhance learning stability. Additionally, [55] apply MoE to expand the capacity of vision-language models, alleviating forgetting in CL. However, these works solely focus on empirical methods, lacking theoretical analysis of how the MoE performs in CL.
@@ 3.1 CL in linear models
General setting. We consider the CL setting with $T$ training rounds. In each round $t\in[T]$, one out of $N$ tasks randomly arrives to be learned by the MoE model with $M$ experts. For each task, we follow most theoretical work on CL by fitting a linear model $f(\mathbf{X})=\mathbf{X}^{\top}\bm{w}$ with ground truth $\bm{w}\in\mathbb{R}^{d}$ (e.g., [12, 34]), which serves as a foundation for understanding DNN generalization performance ([2, 22]). Then, for the task arrival in the $t$-th training round, it corresponds to a linear regression problem, where the training dataset is denoted by $\mathcal{D}_{t}=(\mathbf{X}_{t},\mathbf{y}_{t})$. Here $\mathbf{X}_{t}\in\mathbb{R}^{d\times s_{t}}$ is the feature matrix with $s_{t}$ samples of $d$-dimensional vectors, and $\mathbf{y}_{t}\in\mathbb{R}^{s_{t}}$ is the output vector. In this study, we focus on the overparameterized regime, where $s_{t}<d$. Consequently, there exist numerous linear models that can perfectly fit the data.
Ground truth and dataset. Let $\mathcal{W}=\{\bm{w}_{1},\cdots,\bm{w}_{N}\}$ represent the collection of ground truth vectors of all $N$ tasks. For any two tasks $n,n^{\prime}\in[N]$, we assume $\|\bm{w}_{n}-\bm{w}_{n^{\prime}}\|_{\infty}=\mathcal{O}(\sigma_{0})$, where $\sigma_{0}\in(0,1)$ denotes the variance. Moreover, we assume that task $n$ possesses a unique feature signal $\bm{v}_{n}\in\mathbb{R}^{d}$ with $\|\bm{v}_{n}\|_{\infty}=\mathcal{O}(1)$ ([6, 19]).
@@
In any training round $t\in[T]$, the actual ground truth $\bm{w}_{n_{t}}$ of task arrival $n_{t}$ is unknown. However, according to Definition 1, task $n_{t}$ can be classified into one of $N$ clusters based on its feature signal $\bm{v}_{n_{t}}$. Although the position of $\bm{v}_{n_{t}}$ in feature matrix $\mathbf{X}_{t}$ is not specified for each task $n_{t}$, we can address this binary classification sub-problem over $\mathbf{X}_{t}$ using a single gating network in MoE ([46, 14, 6]). In this context, we aim to investigate whether the MoE model can enhance the learning performance in CL. For ease of exposition, we assume $s_{t}=s$ for all $t\in[T]$ in this paper. Then we will introduce the MoE model in the following subsections.
@@ 3.2 Structure of the MoE model
In each round $t$, as depicted in Figure 1, for task $n_{t}$, the router selects the expert with the maximum gate output $h_{m}(\mathbf{X}_{t},\bm{\theta}_{t}^{(m)})$, denoted as $m_{t}$, from the $M$ experts ([6, 38]). In practice, to encourage exploration across experts and stabilize MoE training, we add perturbations to the router ([46, 14, 6]). Specifically, task $n_{t}$ will be routed to the expert that satisfies
@@
where $r_{t}^{(m)}$ for any $m\in[M]$ is drawn independently from the uniform distribution $\text{Unif}[0,\lambda]$. We analyze in Appendix B that this routing strategy Eq. 1 ensures continuous and stable transitions for tasks. Additionally, the router calculates the softmax gate outputs, derived by
@@ 3.3 Training of the MoE model with key designs
Since we focus on the overparameterized regime, there exist infinitely many solutions that perfectly satisfy $\mathcal{L}_{t}^{tr}(\bm{w}_{t}^{(m_{t})},\mathcal{D}_{t})=0$ in Eq. 3. Among these solutions, gradient descent (GD) starting from the previous expert model $\bm{w}_{t-1}^{(m_{t})}$ at the convergent point provides a unique solution for minimizing $\mathcal{L}_{t}^{tr}(\bm{w}_{t}^{(m_{t})},\mathcal{D}_{t})$ in Eq. 3, which is determined by the following optimization problem ([12, 17, 34]):
@@
Gating network parameters. After obtaining $\bm{w}_{t}^{(m_{t})}$ in Eq. 5, the MoE updates the gating network parameter from $\bm{\Theta}_{t}$ to $\bm{\Theta}_{t+1}$ using GD for the next training round. On one hand, we aim for $\bm{\theta}_{t+1}^{(m)}$ of each expert $m$ to specialize in a specific task, which helps mitigate learning loss caused by the incorrect routing of distinct tasks. On the other hand, the router needs to balance the load among all experts ([14, 46, 31]) to reduce the risk of model overfitting and enhance the learning performance in CL. To achieve this, we introduce our first key design of multi-objective training loss for gating network updates.
Key design I: Multi-objective training loss. First, based on the updated expert model $\bm{w}_{t}^{(m_{t})}$ in Eq. 5, we propose the following locality loss function for updating $\mathbf{\Theta}_{t}$:
@@ 4 Theoretical results on MoE training for CL
In this section, we provide theoretical analysis for the training of expert models and the gating network in Algorithm 1, which further justifies our key designs in Section 3. Specifically, (i) we first support our key design I by proving that the expert model converges fast via updating $\bm{\Theta}_{t}$ under our designed locality loss in Eq. 6. (ii) We then show that our key design II, early termination in updating $\bm{\Theta}_{t}$, is necessary to ensure a stable convergence system state with experts’ balanced loads. For clarity, we study the case with $M>N$ in this section (labeled as $M>N$ version), and further extend the results to the $M<N$ version in appendices. To characterize expert specialization, we first show that each expert’s gate output is determined by the input feature signal $\bm{v}_{n}$ of $\mathbf{X}_{t}$.
@@
The full version containing $M<N$ case and the proof of Lemma 1 are given in Appendix C. According to Lemma 1, the router decides expert $m_{t}$ for task $n_{t}$ based primarily on its feature signal $\bm{v}_{n_{t}}$. Consequently, given $N$ tasks, all experts can be grouped into $N$ sets according to their specialty, i.e., their gating parameter $\bm{\theta}^{(m)}_{t}$’s to identify feature signal $\bm{v}_{n}$, where each expert set is defined as:
@@
The full version and the proof of Proposition 2 are given in Appendix F. According to Proposition 2, if the MoE updates $\bm{\Theta}_{t}$ at any round $t$, the gap between gate outputs of experts within the same expert set converges to $\mathcal{O}(\sigma_{0}^{1.75})$ by round $t_{1}$ (Eq. 12). In contrast, the gap between experts in different sets is sufficiently large, i.e., $\Theta(\sigma_{0}^{0.75})$, indicating that the MoE has successfully diversified experts into different sets at round $t_{1}$ as in Eq. 11. However, unlike MoE in single-task learning that can stop training at any time after the expert models converge (e.g., [4, 31]), MoE in CL requires both the gating network and the experts to be suitably updated with the continuous arrival of new tasks. This is necessary to balance the load on each expert and maximize the system capacity utilization. However, continuing updating $\bm{\Theta}_{t}$ according to Eq. 9 will eventually reduce the output gap between any two experts to $\mathcal{O}(\sigma_{0}^{1.75})$ in Eq. 13 at training round $t_{2}$, causing the router to select wrong experts for subsequent task arrivals and incurring additional training errors.